Chapter 8 Homework Help Save & Exit Submit Check my work 10 Financial analysts forecast Limited Brands (LTD) growth rate for I need help with the whole question and I need the explanation and solution please and thank you.
Answer #1

We know using Grodon growth or constant dividend growth model

Value of stock=Recent Dividend*(1+growth rate)/(required return-growth rate)=0.40*(1+8.5%)/(10.5%-8.5%)=21.70